the perimeter of a playing field is 584 yards. the length is 8 yards shorter than quadruple the width. what are the dimensions?
Answer by srinivas.g(540)

perimeter = 584 yards
length = 4 * width - 8
l = 4 *b -8
but perimeter = 584
2(l+b))= 584
l+b = 292
keep l= 4b -8
4b -8 + b =292
5b-8 =292
5b = 300
b= 60
l = 4*60-8
= 240-8
= 232
length =232
width =60